# Modeling scale-up of particle coating by atomic layer deposition

### v1

@@ III.3 Scale up of continuous processes
Continuous processes differ from batch processes in two ways: first, in a continuous process the precursor is constantly dosed into the reactor volume so that the residence time of the particles inside the reactor, $t_{s}$, and the corresponding normalized residence time, $\tau_{s}$ (Eq. 37), determine the total exposure. Second, the fractional surface coverage of the particles increases as particles move inside the reactor. A key metric for evaluating the manufacturability of continuous processes is therefore the fractional coverage as the particles exit the reactor.
As mentioned in Section III.1, in the case of an ideal self-limited process, the solution for the continuous well-mixed model is identical to that of the well-mixed batch model, except that the dose time, $\tau$, is replaced with the normalized particle residence time, $\tau_{s}$ (Eq. 40). The results shown in Figures 2(A) and 4(A) are therefore applicable to the well-mixed continuous model.
The key features of the plug flow continuous model are shown in Figure 6. In Figure 6(A) we show saturation curves for increasing Damköhler numbers as a function of the normalized particle residence time time, $\tau_{s}$. Here the particle residence time can be viewed as the inverse of the rate or velocity at which particles are being fed into the reactor. As in the case of batch processes, achieving saturation at $\tau_{s}=1$ indicates a transition to the transport-limited regime. In Figure 6(B) we show the fraction of unreacted precursor for the same parameters used in Figure 6(A). In the limit of very high rates of particle insertion ($\tau_{s}\rightarrow 0$), particles move so fast that the precursor reacts with a fully reactive surface. Therefore, the precursor utilization is the highest at very low residence times. High Damköhler numbers lead to almost 100% precursor utilization in this regime. As the residence time increases, the surface coverage of the particles exiting the reactor increases. For high Damköhler numbers, there is a clear transition for $\tau_{s}=1$, when particles leave the reactor fully saturated and the fraction of unreacted precursor starts to increase. Consequently, like in the batch plug flow models, there is a clear signature indicating the optimal particle insertion rate at which saturation is achieved. However, this transition is less abrupt than in batch processes (Fig. 5), since in a continuous process particles are constantly being fed into the reactor and therefore there is still a significant amount of precursor used after reaching saturation.
@@
In Figure 7 we compare the fractional surface coverage achieved at $\tau_{s}=1$ in the well-mixed and plug flow continuous models. As in the case of batch processes, the plug flow model consistently leads to higher fractional surface coverages across the whole range of $\mathrm{Da}$ explored. This highlights the importance of reactor design and precursor delivery to ensure a faster transition to a transport limited regime and an optimal precursor utilization.
@@ III.4 Extension to soft-saturating processes
Figure 8: Saturation curves in a well-mixed batch process for a soft-saturating ALD process. Results are shown for increasing values of the Damköhler number of the main surface reaction pathway: (A) $\mathrm{Da}_{1}=1$; (B) $\mathrm{Da}_{1}=10$; (C)$\mathrm{Da}_{1}=50$. In all cases, $\mathrm{Da}_{2}=0.1\times\mathrm{Da}_{1}$.
@@ IV Discussion
The results obtained also emphasize the importance of reactor design in ensuring a fast and efficient transition from a reaction-limited to a transport-limited regime: the plug flow model consistently leads to faster processes and higher precursor utilization in both batch and continuous ALD configurations. In the batch process case, fluidized bed reactors are one way of achieving this cross flow: indeed, in a prior work we showed good qualitative and quantitative agreement between the plug flow model and experimental results of particle coating by TMA/waterLu et al. (2022). In the continuous case, plug flow configurations show additional properties, such as a self-extinguishing behavior: as shown in Figure 6, the process conditions can be optimized to achieve saturation while minimizing the fraction of unreacted precursor. This can prove beneficial in spatial ALD configurations where isolation between different precursor zones is key.
Finally, the presence of soft-saturating processes shifts the transition to a transport limited regime to higher Damköhler numbers of the leading, more reactive process due to the need to accommodate slower reaction pathways. However, as shown in Figure 8, running processes in a sub-saturating regime can provide a viable tradeoff between throughput and precursor utilization.
@@ V Conclusions
In this work, we have explored the scalability of different strategies for particle coating by atomic layer deposition by analyzing the transition from reaction-limited to transport-limited regime. We have introduced four simple models encompassing both batch and continuous processes of particle coating with agitation. These models depend primarily on inputs available from either knowledge about the ALD process or the experimental conditions. They can help evaluate the scalability of different ALD processes and their implementation at manufacturing scale, something relevant for cost-sensitive applications such as catalysis, energy storage, and decarbonization. The models are available as part of the Python package aldsim. Specific models have also been released as a spreadsheet and made available in the GitHub repository: https://github.com/aldsim/aldsim.

### v2

@@ III.3 Scale up of continuous processes
Continuous processes differ from batch processes in two ways: first, in a continuous process the precursor is constantly dosed into the reactor volume so it is the residence time of the particles inside the reactor, $t_{s}$, what determines the total exposure. Second, the fractional surface coverage of the particles increases as particles move inside the reactor. The relevant observable of a continuous process is therefore the fractional surface coverage as particles exit the reactor. The exit point is defined by the reactor length $L$.
As mentioned in Section III.1, in the case of an ideal self-limited process, the solution for the continuous well-mixed model (Eq. 42) is identical to that of the well-mixed batch model, except that the dose time, $\tau$, is replaced with the normalized particle residence time, $\tau_{s}$ (Eq. 39). The results shown in Figures 2(A) and 4(A) are therefore applicable to the well-mixed continuous model.
The key features of the plug flow continuous model are shown in Figure 6. In Figure 6(A) we show saturation curves for increasing Damköhler numbers as a function of the normalized particle residence time, $\tau_{s}$. The normalized particle residence time is inversely proportional to the velocity at which the particles are being moved into the reactor, and proportional to the reactor size (Eq. 39). When particles are inserted in the reactor at a very high rate ($\tau_{s}\rightarrow 0$), particles leave the reactor largely uncoated. $\tau_{s}=1$ corresponds to the case where the number of surface sites and the number of precursor molecules inserted in the reactor per unit time are the same. Consequently, as in the case of batch processes, a linear increase of surface coverage with residence time that leads to saturation at $\tau_{s}=1$ indicates that the system is in the transport-limited regime.
In Figure 6(B) we show the fraction of unreacted precursor for the same parameters used in Figure 6(A). At very low residence times, the fraction of precursor leaving the reactor is very small, indicating almost 100% precursor utilization. For $\tau_{s}>1$, the rate at which precursor is fed into the reactor exceeds that of the surface sites, and the fraction of unreacted precursor starts to increase. Both Figs 6(A) and 6(B) show that, at high Damköhler numbers, there is a clear transition at $\tau_{s}=1$ separating the undersaturated regime, where growth is still not fully saturated and precursor is fully utilized, and the saturated regime, where the surface fractional coverage is very high and precursor utilization starts to decrease. Consequently, like in the batch plug flow case, techniques such as downstream mass spectrometry should provide a clear signature indicating the optimal particle insertion rate at which saturation is achieved. However, this transition is less abrupt than in batch processes (Fig. 5).
@@
In Figure 7 we compare the fractional surface coverage achieved at $\tau_{s}=1$ in the well-mixed and plug flow continuous models. As in the case of batch processes, continuous processes where precursor transport can be approximated with a plug flow model consistently lead to higher fractional surface coverages across the whole range of $\mathrm{Da}$ explored. This highlights the importance of reactor design and precursor delivery to ensure a faster transition to a transport limited regime and an optimal precursor utilization.
@@ III.4 Extension to soft-saturating processes
Figure 8: Saturation curves in a well-mixed batch process for a soft-saturating self-limited surface kinetics. Results are shown for increasing values of the Damköhler number of the main surface reaction pathway: (A) $\mathrm{Da}_{1}=1$; (B) $\mathrm{Da}_{1}=10$; (C)$\mathrm{Da}_{1}=50$. In all cases, $\mathrm{Da}_{2}=0.1\times\mathrm{Da}_{1}$.
@@ IV Discussion
The results obtained also emphasize the importance of reactor design in ensuring a fast and efficient transition from a reaction-limited to a transport-limited regime: the plug flow model consistently leads to faster processes and higher precursor utilization in both batch and continuous ALD configurations. In the batch process case, fluidized bed reactors are one way of achieving this cross flow. In the continuous case, plug flow configurations show additional properties, such as a self-extinguishing behavior: as shown in Figure 6, the process conditions can be optimized to achieve saturation while minimizing the fraction of unreacted precursor. This can prove beneficial in spatial ALD configurations where isolation between different precursor zones is key. We have also observed that, in the case of soft-saturating processes where there is a secondary slower reaction pathway, the transition from reaction to a transport limited regime shifts to higher Damköhler numbers. However, as shown in Figure 8, running processes in a sub-saturating regime can provide a viable tradeoff between throughput and precursor utilization.
Finally, there is the question of how well these models agree with experiments. As discussed in Section III.2, for the case of fluidized reactors we have observed good qualitative and quantitative agreement between the plug flow model and experimental results of particle coating by TMA/waterLu et al. (2022). The punch-through of the precursor signal as measured by mass spectrometry, the near 100% precursor consumption prior to reaching full saturation, and even the predicted saturation dose times were found to agree well with the model predictions. However, further experimental work is needed in order to validate the approximations in other promising reactor configurations. The models explored in this work also don’t consider additional sources of non-idealities that can affect process scale up. One example is particle agglomeration, where the need for the precursor to diffuse within the agglomerates can lead to longer saturation times. These effects require more detailed simulation approaches that are beyond the scope of this workGrillo, Kreutzer, and van Ommen (2015); Jin, Kleijn, and van Ommen (2017).
@@ V Conclusions
In this work, we have explored the scalability of different strategies for particle coating by atomic layer deposition by analyzing the transition from reaction-limited to transport-limited regime. We have introduced four simple models encompassing both batch and continuous processes of particle coating with agitation. These models depend primarily on inputs available from either knowledge about the ALD process or the experimental conditions. They can help evaluate the scalability of different ALD processes and their implementation at manufacturing scale, something relevant for cost-sensitive applications such as catalysis, energy storage, and decarbonization. The models are available as part of the Python package aldsim, and made available in the GitHub repository: https://github.com/aldsim/aldsim.
@@ Acknowledgements.
Author Declarations
Conflict of interest
The authors have no conflicts to disclose
@@
Appendix A Non-dimensional equations
This document includes the non-dimensional equations for the eight models considered in the manuscript.
A.1 Batch process, well-mixed
Particles are homogeneously mixed, and precursor transport is modeled using the well-mixed approximation (Section II.2.1).
A.1.1 Single reaction pathway
$(1-x)$ $=$ $\mathrm{Da}(1-\Theta)x$ (45)
$\frac{d\Theta}{d\tau}$ $=$ $\mathrm{Da}(1-\Theta)x$ (46)
where:
$x$ $=$ $n/n_{0}$ (47)
$\tau$ $=$ $t/t_{0}$ (48)
$\mathrm{Da}$ $=$ $\frac{S}{V}f\beta_{0}\frac{1}{4}v_{th}t_{\mathrm{res}}$ (49)
$t_{0}$ $=$ $t_{\mathrm{res}}\frac{S}{s_{0}n_{0}V}$ (50)
A.1.2 Soft-saturating process
$\begin{split}(1-x)=&\left[(1-f)\mathrm{Da}_{a}(1-\Theta_{a})\right.\\ &\left.+f\mathrm{Da}_{b}(1-\Theta_{b})\right]x\end{split}$ (51)
and
$\frac{d\Theta_{a}}{d\tau}$ $=$ $\mathrm{Da}_{a}(1-\Theta_{b})x$ (52)
$\frac{d\Theta_{b}}{d\tau}$ $=$ $\mathrm{Da}_{b}(1-\Theta_{b})x$ (53)
where $x$, $\tau$ and $t_{0}$ are defined in Eqs. 47, 48, and 50 and:
$\mathrm{Da}_{a}$ $=$ $\frac{S}{V}\beta_{a}\frac{1}{4}v_{th}t_{\mathrm{res}}$ (54)
$\mathrm{Da}_{b}$ $=$ $\frac{S}{V}\beta_{b}\frac{1}{4}v_{th}t_{\mathrm{res}}$ (55)
A.2 Batch process, plug flow
Particles are homogeneously mixed, and the precursor transport is modeled using a plug flow model. This corresponds to situations where the precursor is forced to flow through volatilized particles (Section II.2.2).
A.2.1 Single reaction pathway
$\frac{dx}{d\xi}$ $=$ $-\mathrm{Da}(1-\Theta)x$ (56)
$\frac{d\Theta}{d\tau}$ $=$ $\mathrm{Da}(1-\Theta)\bar{x}$ (57)
[…]
where $x$, $\tau$, $\mathrm{Da}$, and $t_{0}$ are defined in Eqs. 47, 48, 49, and 50 with
$t_{\mathrm{res}}=\frac{L}{u}$ (59)
and
$\xi=\frac{z}{L}$ (60)
A.2.2 Soft-saturating process
$\begin{split}\frac{dx}{d\xi}=&-\left[(1-f)\mathrm{Da}_{a}(1-\Theta_{a})\right.\\ &+\left.f\mathrm{Da}_{b}(1-\Theta_{b})\right]x\end{split}$ (61)
and
$\frac{d\Theta_{a}}{d\tau}$ $=$ $\mathrm{Da}_{a}(1-\Theta_{b})\bar{x}$ (62)
$\frac{d\Theta_{b}}{d\tau}$ $=$ $\mathrm{Da}_{b}(1-\Theta_{b})\bar{x}$ (63)
[…]
with $\mathrm{Da}_{a}$ and $\mathrm{Da}_{b}$ defined in Eqs. 54 and 55.
A.3 Continuous process, well-mixed
Particles move along the reactor and homogeneously mixed in the plane perpendicular to the direction of movement and stratified in the direction of movement. Precursor transport is modeled using the well-mixed approximation (Section II.2.3).
A.3.1 Single reaction pathway
$(1-x)$ $=$ $\mathrm{Da}(1-\Theta)x$ (65)
$\frac{d\Theta}{d\xi}$ $=$ $\tau_{s}\mathrm{Da}(1-\Theta)x$ (66)
[…]
Here:
$\tau_{s}=\frac{L}{vt_{0}}$ (68)
and $x$, $\tau$, $\mathrm{Da}$, $t_{0}$ and $\xi$ are defined in Eqs. 47, 48, 49, and 50, and 60.
A.3.2 Soft-saturating process
$\begin{split}(1-x)&=\left[(1-f)\mathrm{Da}_{a}(1-\bar{\Theta}_{a})\right.\\ &\left.+f\mathrm{Da}_{b}(1-\bar{\Theta}_{b})\right]x\end{split}$ (69)
and
$\frac{d\Theta_{a}}{d\xi}$ $=$ $\tau_{s}\mathrm{Da}_{a}(1-\Theta_{b})x$ (70)
$\frac{d\Theta_{b}}{d\xi}$ $=$ $\tau_{s}\mathrm{Da}_{b}(1-\Theta_{b})x$ (71)
with
[…]
A.4 Continuous process, plug flow
Particles move along the reactor through a continuous feed system. Particles are homogeneously mixed in the plane perpendicular to the direction of movement and stratified in the direction of movement. Precursor moves downstream along the direction of the particles and its transport is modeled using a plug flow approximation (Section II.2.4).
A.4.1 Single reaction pathway
$\frac{dx}{d\xi}$ $=$ $-\mathrm{Da}(1-\Theta)x$ (74)
$\frac{d\Theta}{d\xi}$ $=$ $\tau_{s}\mathrm{Da}(1-\Theta)x$ (75)
A.4.2 Soft-saturating process
$\begin{split}\frac{dx}{d\xi}=&-\left[(1-f)\mathrm{Da}_{a}(1-\Theta_{a})\right.\\ &\left.+f\mathrm{Da}_{b}(1-\Theta_{b})\right]x\end{split}$ (76)
and
$\frac{d\Theta_{a}}{d\xi}$ $=$ $\tau_{s}\mathrm{Da}_{a}(1-\Theta_{b})x$ (77)
$\frac{d\Theta_{b}}{d\xi}$ $=$ $\tau_{s}\mathrm{Da}_{b}(1-\Theta_{b})x$ (78)